OpenStudy (anonymous):

find all roots of this equation x^3-5x^2+2x+8=0

OpenStudy (anonymous):

x =2, -1, 4

OpenStudy (anonymous):

how did u do that though

OpenStudy (anonymous):

for the 1st root you hav to find it just by substituting a satisfactory value..i.e 2. it's not hard if you hv got a proper practice. then, x^3-5x^2+2x+8 = (x-2)*(x^2 + bx -4) you ll get b = -3 , i.e (x^2-3x-4) = 0 ==>> (x-4)(x+1) = 0

OpenStudy (anonymous):

super bad at math thank you so much for your help... i kind of get it!